# Observing network dynamics through sentinel nodes

### v1

@@ Acknowledgements
We analyzed the obtained $\ln\varepsilon$ values by a multi-way ANOVA as in section S4. Even without knowledge of the test dynamics, our model explains a substantial portion of the variance in $\ln\varepsilon$ ($R^{2}=0.58$), although it is less than when we do know the test dynamics (see section S4). Each of the four independent variables is highly significant (training dynamics: $df=3$, $F=2980.2$, $p<10^{-7}$; test dynamics: $df=3$, $F=24920.9$, $p<10^{-7}$; network: $df=8$, $F=643.8$, $p<10^{-7}$; node set type: $df=6$, $F=2643.2$, $p<10^{-7}$). Note that the large number (i.e., $75,600$) of observations in part explains the small $p$-values.
@@
Table S3 shows the differences in average $\ln\varepsilon$, estimated by Tukey’s honestly significant difference test. Despite not knowing the test dynamics, optimized node sets are still associated with the smallest $\varepsilon$, followed by the degree-preserving, heuristic algorithms, and completely random node sets. Although the difference between several pairs of the four heuristic algorithms in terms of $\ln\varepsilon$ is statistically significant, the magnitude of the differences among them is small.
@@
Table S3: Differences in average $\ln\varepsilon$ between different node set types when we do not know the test dynamics. The differences are computed by a Tukey’s honestly significant difference test. The 95% confidence intervals (CI) of the differences and associated $p$-values, adjusted for multiple comparisons, are also shown.
@@
Optimized -- Random $-1.748$ $[-1.804,-1.692]$ $<10^{-7}$
Degree-preserving -- Random $-1.310$ $[-1.366,-1.254]$ $<10^{-7}$
$k$-constrained -- Random $-0.170$ $[-0.226,-0.114]$ $<10^{-7}$
$k$-quantiled -- Random $-0.138$ $[-0.194,-0.082]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Random $-0.262$ $[-0.319,-0.206]$ $<10^{-7}$
Community-based -- Random $-0.157$ $[-0.213,-0.101]$ $<10^{-7}$
Degree-preserving -- Optimized $0.438$ $[0.381,0.494]$ $<10^{-7}$
$k$-constrained -- Optimized $1.578$ $[1.522,1.634]$ $<10^{-7}$
$k$-quantiled -- Optimized $1.610$ $[1.554,1.666]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Optimized $1.485$ $[1.429,1.541]$ $<10^{-7}$
Community-based -- Optimized $1.591$ $[1.535,1.647]$ $<10^{-7}$
$k$-constrained -- Degree-preserving $1.140$ $[1.084,1.196]$ $<10^{-7}$
$k$-quantiled -- Degree-preserving $1.173$ $[1.117,1.229]$ $<10^{-7}$
$k_{\rm nn}$-constrained -- Degree-preserving $1.048$ $[0.992,1.104]$ $<10^{-7}$
Community-based -- Degree-preserving $1.153$ $[1.097,1.209]$ $<10^{-7}$
$k$-quantiled -- $k$-constrained $0.032$ $[-0.024,0.088]$ $0.613$
$k_{\rm nn}$-constrained -- $k$-constrained $-0.092$ $[-0.149,-0.036]$ $2.37\times 10^{-5}$
Community-based -- $k$-constrained $0.013$ $[-0.043,0.069]$ $0.993$
$k_{\rm nn}$-constrained -- $k$-quantiled $-0.125$ $[-0.181,-0.069]$ $<10^{-7}$
Community-based -- $k$-quantiled $-0.019$ $[-0.075,0.037]$ $0.950$
Community-based -- $k_{\rm nn}$-constrained $0.106$ $[0.049,0.162]$ $6\times 10^{-7}$
@@
Figure S14: Approximation of the average state, $\overline{x}$, of the coupled double-well dynamics on the dolphin network when we also optimize node weights. (a) $n=1$. (b) $n=2$. (c) $n=3$. (d) $n=4$. The tan lines represent the weight-optimized approximation. The purple lines represent the corresponding approximation when we do not optimize node weights.
S8.2 Effect of node set type on approximation error
To quantify performances of the optimized node sets in which we also optimize the node weights, we built an ANOVA model using $\ln\varepsilon$ as the dependent variable, as we did in section S4. We generated 100 weight-optimized node sets for each pair of dynamics and network. For comparison, we use the same completely random and combinatorially optimized node sets (i.e., without optimizing node weights) from section S4. The three independent variables were dynamics, network, and node set type (random, optimized, and weight-optimized). All three independent variables were significant (dynamics: $df=3$, $F=5436.06$, $p<10^{-7}$; network: $df=8$, $F=155.76$, $p<10^{-7}$; node set type: $df=2$, $F=12362.80$, $p<10^{-7}$).
We computed the differences between the three node set types with Tukey’s honestly significant difference test. We show the results in Table S4. As described in the main text, the average $\ln\varepsilon$ for the weight-optimized node sets is small, i.e., $46.161/e^{-3.832}=$ 46.16 times smaller than the optimized node sets without the additional weight optimization step.
@@
Optimized -- Random $-5.625$ $[-5.767,-5.483]$ $<10^{-7}$
Weight-optimized -- Random $-9.456$ $[-9.598,-9.315]$ $<10^{-7}$
Weight-optimized -- Optimized $-3.832$ $[-3.973,-3.690]$ $<10^{-7}$
S8.3 Transfer learning
@@
All of the independent variables of the constructed ANOVA model are significant (training dynamics: $df=3$, $F=1353.61$, $p<10^{-7}$; test dynamics: $df=3$, $F=11278.18$, $p<10^{-7}$; network: $df=8$, $F=285.22$, $p<10^{-7}$; node set type: $df=2$, $F=5171.33$, $p<10^{-7}$), and the ANOVA model fits the data well ($R^{2}=0.61$). As reported in the main text, there is no significant difference between weight-optimized node sets and our standard optimized node sets (i.e., without node-weight optimization) when we optimize on a dynamics that is not the test dynamics ($p=0.591$).

### v3

@@ C Heuristic algorithms for node set selection
As in section S4, we analyzed $\ln\varepsilon$ as the dependent variable in a multi-way ANOVA. We excluded data from the ER network for the same reason as that stated in section S4. Even without knowledge of the test dynamics, our model explains a substantial portion of the variance in $\ln\varepsilon$ ($R^{2}=0.64$), although it is less than when we do know the test dynamics (see section S4). Each of the four independent variables is highly significant (training dynamics: $df=3$, $F=2,355$, $p<10^{-7}$; test dynamics: $df=3$, $F=27,863$, $p<10^{-7}$; network: $df=18$, $F=605$, $p<10^{-7}$; node set type: $df=2$, $F=9,561$, $p<10^{-7}$). Note that the large number (i.e., $68,400$) of observations in part explains the small $p$ values.
@@
Table S19 shows the differences in average $\ln\varepsilon$, estimated by Tukey’s honestly significant difference test. Despite not knowing the test dynamics, optimized node sets are still associated with the smallest $\varepsilon$, followed by the degree-preserving node sets and then by the completely random node sets.
@@
Table S19: Differences in average $\ln\varepsilon$ between different node set types when we do not know the test dynamics. The differences are computed by a Tukey’s honestly significant difference test. The 95% confidence intervals (CI) of the differences and associated $p$ values, adjusted for multiple comparisons, are also shown.
@@
Optimized - Random 2.006-2.006 [2.042,1.970][-2.042,-1.970] $10^{-7}$
Degree-preserving - Random 1.587-1.587 [1.623,1.551][-1.623,-1.551] $10^{-7}$
Degree-preserving - Optimized $0.419$ $[0.383,0.455]$ $10^{-7}$
Despite these statistical results, an analysis of the Prosper network, which is the largest network used in the present study, reveals a different picture. On average, Fig. S25 indicates that the degree-preserving node sets outperform the optimized node sets on the test dynamics in six out of the twelve pairs of training and test dynamics (i.e., the large triangle is below the large circle in six panels in Fig. S25a). This result contrasts to the case of the dolphin network shown in Fig. 5a in the main text, in which the degree-preserving node sets outperform the optimized node sets on the test dynamics only in two pairs of training and test dynamics. A major difference between the two networks used is the size of the network. Therefore, to explore generality of this observation, for each of the 19 networks, we counted the number of pairs of training and test dynamics for which the degree-preserving node sets perform better than the optimized node sets on average. We show the results as a function of the number of nodes, $N$, in Fig. S28a. The figure suggests that the degree-preserving node sets perform relatively better for larger $N$. The Pearson’s correlation coefficient between $\log N$ and the count is $r=0.52$ with $p=0.021$.
We also found that the degree-preserving node sets tend to perform better when either the training or test dynamics is the gene-regulatory dynamics. To quantify this observation, for each pair of training and test dynamics, we counted the number of networks for which the degree-preserving node sets outperform the optimized node sets. We show the results in Fig. S28b.
Figure S28: The number of cases in which the degree-preserving node sets outperform the optimized node sets in transfer learning. (a) Count of pairs of training and test dynamics for which the degree-preserving node sets are better. A circle represents a network. (b) Count of networks for which the degree-preserving node sets are better, given the pair of training and test dynamics. In both (a) and (b), we excluded the ER network because it has limited heterogeneity in the degree distribution.
Appendix S13 Incomplete observation of the network
To investigate the performance of optimized sentinel node sets when some nodes or edges are not observable, we carry out the following stress test. For a given network, we first remove a given fraction of edges, $f$, selected uniformly at random. Then, we run our optimization algorithm on the reduced network to determine a sentinel node set. Then, we test its performance on the original network. This scenario corresponds to the case in which some edges are unobservable but contribute to the ongoing network dynamics. We use the coupled double-well dynamics for demonstration.
We show in Fig. S29a results when some edges in the Proximity network are missing. For each $f$ value used in the figure, we generate one reduced network and run the optimization algorithm five times per reduced network. A circle in the figure corresponds to a node set and shows the approximation error measured against $\{x_{1}^{*},\ldots,x_{N}^{*}\}$ for the dynamics run on the original network. The blue and green symbols represents the approximation error for node sets optimized for the original network and that for completely random node sets, respectively, for reference. By noting the logarithmic scale on the vertical axis, we suggest that the approximation error remains overall smaller than that for the completely random node sets roughly when $f\leq 0.3$.
We also run similar experiments removing a fraction $f$ of nodes, rather than edges, uniformly selected at random. In this scenario, we also remove all the edges incident to the removed nodes. We show the approximation error for the node set optimized for reduced networks, where the approximation error is measured for the dynamics run on the original network, in Fig. S29b. Qualitatively similar to the case of edge removal, the approximation error remains smaller than the case of completely random node sets roughly when $f\leq 0.15$.
Results for a larger network are similar (see Fig. S29c and S29d), whereas the approximation error grows faster as we remove edges in this network (see Fig. S29c) compared to the first network (see Fig. S29a). We conclude that our framework for selecting good sentinel node sets is tolerant to missing observation of up to a nonnegligible fraction (such as $f=0.1$ or $0.2$) of nodes or edges.
Figure S29: Sentinel node approximation under incomplete observation of networks. (a) Proximity network under edge removal. (b) Proximity network under node removal. (c) Email network under edge removal. (d) Email network under node removal. We use the coupled double-well dynamics and show the approximation error, $\varepsilon$, with $n=\lfloor\ln N\rfloor$ sentinel nodes. Each marker in the main plot represents the approximation error, $\varepsilon$, for a sentinel node set optimized on the reduced network in which some edges or nodes are unobserved. The obtained sentinel node set is tested against the dynamics on the original network. For each network, we also show the results for $100$ optimized node sets (blue circles) and 100 completely random node sets (green squares) for the original network. We show $\varepsilon$ for the node sets using the original network between panels (a) and (b) and between panels (c) and (d) to ease the comparison between the reduced and original networks.
Appendix S14 When the coupling strength of only some edges vary
In the main text, we assumed that the coupling strength between nodes, $D$, is globally shared and tuned. In this section, we investigate the extent to which the performance of optimized sentinel node sets is robust when this assumption is violated. We assume that the coupling strength for a fraction $f$ of edges remains fixed at the initial value of $D$ and that the $D$ value for the other edges varies as it does in the analysis in the main text. If $f$ is small, then the dynamics as we vary the $D$ value would be similar to the original case, i.e., when all edges share the varying $D$ value. We computed sentinel node sets when $D$ is globally tuned (i.e., $f=0$) and tested the performance of the obtained sentinel node sets against simulations with $f>0$. We considered five methods to select a fraction $f$ of edges for which the coupling strength is fixed over the course of simulations. We select the fraction $f$ of edges with (i) the lowest edge betweenness centrality, (ii) highest edge betweenness centrality, (iii) lowest degree sum, (iv) highest degree sum, or (iv) uniformly at random. The degree sum of the edge is defined as the sum of the degree of two nodes forming the edge.
We show the approximation error for combinations of $f=\{0.01,0.05,0.1,0.2,0.3,0.4,0.5\}$ and one of the five edge selection methods for the Proximity network in Fig. S30a. We used the coupled double-well dynamics. For $100$ optimized node sets determined under the globally tuned coupling strength, which is used in all other parts of this study, the approximation error under the same globally tuned coupling strength is shown by the blue circles. Each row of the $100$ circles of the same darkness of gray circles corresponds to a value of $f$. A darker gray corresponds to a smaller value of $f$. For example, the black circles in the first of the seven rows of circles for each edge selection method correspond to $f=0.01$. The lightest gray circles in the seventh row correspond to $f=0.5$. The $100$ gray circles in each row correspond to the same $100$ optimized node sets shown in blue, but evaluated for the scenario in which a fraction $f$ of edges has a fixed coupling strength while the other fraction 1f1-f of edges has a varying coupling strength. The green squares show the approximation error for $100$ completely random node sets in the case of the globally shared $D$ as reference.
Figure S30a suggests that the approximation error when some edges stick to its initial coupling strength remains smaller than the completely random case when $f\leq 0.1$ overall for four of the five edge selection methods (i.e., except for “lowest degree sum”). Note that the horizontal axis is on the logarithmic scale. For the “lowest degree sum”, the approximation error is smaller than the completely random node sets only for $f\leq 0.05$. The results are similar for a larger network, as shown in Fig. S30b. We consider that optimized node sets determined with $f=0$ is transferrable only up to a relatively small value of $f$ because selectively changing edges with the highest/lowest betweenness or degree sum distorts the effective weighted network considerably even if $f$ is small. In fact, when a fraction $f$ of edges is uniformly randomly selected, a reasonable quality of transferability persists up to $f=0.4$ for both networks (see “Random” in Fig. S30). Therefore, we conclude that our sentinel node selection method bears some robustness (i.e., up to $f=0.05$ to $0.1$ even when particular edges are chosen to have a fixed coupling strength and up to $f=0.4$ when edge selection is uniformly at random) under this scenario.
Figure S30: Sentinel node approximation when the coupling strength gradually varies only for a fraction 1f1-f of edges. (a) Proximity network. (b) Email network. Each gray and blue circle represents a sentinel node set optimized for the original network. The darker gray circles correspond to a larger value of $f$. The green squares represent completely random node sets. The blue and green symbols show the approximation error, $\varepsilon$, assessed for the original case, i.e., when the coupling strength is shared by all edges. The gray symbols show $\varepsilon$ assessed for the case in which a fraction $f$ of edges have a fixed coupling strength, and the remaining edges have a shared varying coupling strength. We use the coupled double-well dynamics and set $n=\lfloor\ln N\rfloor$.
Appendix S15 Sentinel node approximation for brain data
To demonstrate transfer learning with multivariate time series data for which the dynamics generating the data is unknown, we used the resting-state functional magnetic resonance imaging (fMRI) data recorded from healthy human individuals collected in the Human Connectome Project (HCP) [25]. The participants are $1,200$ adults between 22223535 years old who underwent four sessions of 15-min EPI sequence with a 3T Siemens Connectome-Skyra (TR =0.72=0.72 s, TE =33.1=33.1 ms, $72$ slices, $2.0$ mm isotropic, field of view (FOV) $208\times 180$ mm) and a T1-weighted sequence (TR =2.4=2.4 s, TE =2.14=2.14 ms, $0.7$ mm isotropic, FOV $224\times 224$ mm). The resting-state fMRI data of each participant are composed of two sessions, and each of the two sessions is broken down into a Left-Right (LR) and Right-Left (RL) phases. Each phase provides a time series with 1,200 time points. We used the preprocessed time series data provided by the HCP, called the “node timeseries (individual subjects)” data. These data are from 1,003 participants having complete resting-state fMRI data over four phases. The HCP provides the data from the four phases as a concatenated time series of 4,800 time points. Each node is an ICA (independent component analysis) component as a result of parcellation of the brain. We used the time series with the largest number of nodes, i.e., $N=300$.
Some participants have nodes showing outlier behavior. To remove the data from such participants, for each participant, we first computed $\overline{x}_{i}$, $\forall i$ for the test set, i.e., the last 100 time points. We then computed the mean and standard deviation of the 100 values of $\overline{x}_{i}$, denoted by $\mu$ and $\sigma$, respectively. If any $\left|x_{i}\right|$ at any time point is larger than $\mu+5\sigma$, we removed that participant’s data from further consideration. Using this rule, we removed $103$ out of $1,003$ participants, leaving $900$ participants.
For each of the $900$ participants, we estimated a network by the Pearson correlation coefficient between each pair of nodes. We used the first 4,700 time points from each node to calculate the correlation. Although the Pearson correlation is a naive method, various methods for estimating networks from correlation data have pros and cons [108]. Because our goal of the data analysis is to demonstrate transfer learning in the setting where the dynamics producing time series data is unavailable, we resorted to a simplest network estimation method, i.e., Pearson correlation. We discarded negative correlation coefficient values to obtain a (positively) weighted network. Although the “node timeseries (individual subjects)” data set contains two networks per participant, they are correlation-based. We computed the correlation network from the time series data on our own because, in this manner, we can separate the training data (i.e., the first 4,700 time points) to estimate the network and the test data (i.e., the last 100 time points) for running our optimization algorithm and assessing the performance of the obtained sentinel node sets.
We ran the coupled double-well dynamics on the network estimated for each participant for the range of $D$ values used in the main text. Then, we ran the optimization algorithm to identify a set of $n=\lfloor\ln N\rfloor=5$ sentinel nodes. We removed $\sum_{\ell=1}^{L}\overline{x}_{\ell}$ in the definition of $\varepsilon$, i.e., Eq. (3) in the main text, because the present $x_{i}$ can take both positive and negative values and this sum only contributes a normalization factor without affecting the optimization. For each participant, we repeated this procedure $100$ times, and computed $\varepsilon$ for each optimized node set. We also generated the same number of degree-preserving and completely random node sets and calculated $\varepsilon$ for each node set. We visually compare the average $\varepsilon$ between the node set types across the $900$ participants in the main text. We find that, although the tendency is weak, optimal node sets tend to yield a smaller $\varepsilon$ than completely random node sets. We used a mixed effects model in which the dependent variable was approximation error, the fixed independent variable was node set type (reference: completely random node sets), and there was a random intercept for each individual. Because $\varepsilon$ is not widely distributed, we used $\varepsilon$, not $\ln\varepsilon$, as the dependent variable. We found that $\varepsilon$ for the optimal node sets was 5.81% smaller than that for the completely random node sets (see Table S20).
To assess the generality of these results with respect to the choice of the dynamics, we ran the same analysis for the mutualistic species, SIS, and gene-regulatory dynamics. In addition, we ran the same analysis for the coupled Wilson-Cowan model for neuronal dynamics given by
$\frac{\text{d}x_{i}}{\text{d}t}=-x_{i}+\sum_{j=1}^{N}\frac{A_{ij}}{1+e^{\mu-\tau x_{j}}},$ (S15)
where $x_{i}$ represents the activation level of a mass of neurons, $\mu$ controls the activation threshold, and $\tau$ controls the slope of the activation function [10]. We set $\mu=3$ and $\tau=1$ following [10]. We show the results in Table S20. The table indicates that the result that optimal and degree-preserving node sets, in particular the former, provide a better approximation to $\overline{x}$ than completely random node sets remains the same for the different dynamics.
In Table S20, we analyzed each dynamics separately. We further confirmed these results by evaluating the same statistical model as the one used in Table S20 (i.e., mixed effects with random intercepts for participants, etc.) for all five dynamics combined, except that we added the model of dynamics used as the training dynamics as another categorical independent variable. The results of this analysis, shown in Table S21, are similar to those of analyzing the dynamics separately, shown in Table S20. In particular, the training dynamics used to generate node sets does not have a strong effect on the approximation error when compared to the type of node set. Over all dynamics, optimized node sets on average achieve 6.40% smaller approximation error than completely random node sets; degree-preserving node sets achieve 3.13% smaller approximation error.
Table S20: Improvement, with respect to completely random node sets, of optimized and degree-preserving node sets on the fMRI data for each of five training dynamics. The coefficient estimates $b$, as well as the associated $t$ and $p$ values, come from separate mixed-effects models for each training dynamics.
Training dynamics Node set type % improvement $b$ $t$ $p$
Coupled double-well Optimized $5.81$ 103.7-103.7 7.727-7.727 $10^{-7}$
Coupled double-well Degree-preserving $3.33$ 59.48-59.48 4.431-4.431 $9.96\times 10^{-6}$
Mutualistic species Optimized $5.95$ 106.7-106.7 8.429-8.429 $10^{-7}$
Mutualistic species Degree-preserving $1.72$ 30.87-30.87 2.440-2.440 0.0148
SIS Optimized $3.94$ 70.43-70.43 5.656-5.656 $10^{-7}$
SIS Degree-preserving $3.96$ 70.74-70.74 5.680-5.680 $10^{-7}$
Gene-regulatory Optimized $7.79$ 139.1-139.1 12.23-12.23 $10^{-7}$
Gene-regulatory Degree-preserving $3.45$ 61.58-61.58 5.413-5.413 $10^{-7}$
Wilson-Cowan Optimized $8.5$ 152.2-152.2 11.95-11.95 $10^{-7}$
Wilson-Cowan Degree-preserving $3.18$ 56.81-56.81 4.461-4.461 $8.68\times 10^{-6}$
Table S21: Statistical results for a single mixed effects model, including all five dynamics, for the fMRI data. The dependent variable is the approximation error. There were two categorical fixed effects: training dynamics (reference: coupled double-well dynamics) and node set type (reference: completely random node sets). There was a random effect for participant.
$b$ $t$ $p$
Intercept $1787$ $146.5$ $10^{-7}$
Training dynamics
   Mutualistic species $15.14$ $2.108$ 0.0350
   SIS $8.727$ $1.215$ 0.2244
   Gene-regulatory 10.94-10.94 1.523-1.523 0.1277
   Wilson-Cowan 10.84-10.84 1.509-1.509 0.1313
Node set type
   Optimized 114.4-114.4 20.57-20.57 $10^{-7}$
   Degree-preserving 55.90-55.90 10.05-10.05 $10^{-7}$
@@ A An example
Figure S31: Approximation of the average state, $\overline{x}$, of the coupled double-well dynamics on the dolphin network when we also optimize node weights. (a) $n=1$. (b) $n=2$. (c) $n=3$. (d) $n=4$. The tan lines represent the weight-optimized approximation. The purple lines represent the corresponding approximation when we do not optimize node weights. In (a), the tan line completely covers the purple line.
B Effect of node set type on approximation error
To quantify performances of the optimized node sets in which we also optimize the node weights, we built an ANOVA model using $\ln\varepsilon$ as the dependent variable, as we did in section S11C. We generated 100 weight-optimized node sets for each pair of dynamics and network. For comparison, we use the same completely random and combinatorially optimized node sets (i.e., without optimizing node weights) from section S11C. The three independent variables were dynamics, network, and node set type (random, optimized, and weight-optimized). All three independent variables were significant (dynamics: $df=3$, $F=5,436.06$, $p<10^{-7}$; network: $df=8$, $F=155.76$, $p<10^{-7}$; node set type: $df=2$, $F=12,362.80$, $p<10^{-7}$).
We computed the differences between the three node set types with Tukey’s honestly significant difference test. We show the results in Table S22. As described in the main text, the average $\ln\varepsilon$ for the weight-optimized node sets is small, i.e., $1/e^{-3.832}=46.16$ times smaller than the optimized node sets without the additional weight optimization step.
@@ B Effect of node set type on approximation error
Optimized - Random 5.625-5.625 [5.767,5.483][-5.767,-5.483] $10^{-7}$
Weight-optimized - Random 9.456-9.456 [9.598,9.315][-9.598,-9.315] $10^{-7}$
Weight-optimized - Optimized 3.832-3.832 [3.973,3.690][-3.973,-3.690] $10^{-7}$
C Transfer learning
@@
All of the independent variables of the constructed ANOVA model are significant (training dynamics: $df=3$, $F=1,353.61$, $p<10^{-7}$; test dynamics: $df=3$, $F=11,278.18$, $p<10^{-7}$; network: $df=8$, $F=285.22$, $p<10^{-7}$; node set type: $df=2$, $F=5,171.33$, $p<10^{-7}$), and the ANOVA model fits the data well ($R^{2}=0.61$). As reported in the main text, there is no significant difference between weight-optimized node sets and our standard optimized node sets (i.e., without node-weight optimization) when we optimize on a dynamics that is not the test dynamics ($p=0.591$; see Table S23).